Triangle ABC is isosceles with AB=AC, AB=4x-2, BC=4x-3, AC=2x+12 find BC)
Found 2 solutions by Edwin McCravy, stanbon:
Answer by Edwin McCravy(20060)

 
AB = AC 
AB = 4x-2
AC = 2x+12

Substitute 4x-2 for AB and 2x+12 for AC in

           AB = AC
         4x-2 = 2x+12

Add 2 to both sides

           4x = 2x+14

Subtract 2x from both sides:

           2x = 14

Divide both sides by 2

            x = 7

Substitute 7 for x in

           BC = 4x-3
           BC = 4(7)-3
           BC = 28-3
           BC = 25  

Edwin

Answer by stanbon(75887)

triangle abc is isosceles with ab=ac, ab=4x-2, bc=4x-3, ac=2x+12 find bc)
------------------
ab = ac
4x-2 = 2x+12
2x = 14
x = 7
-----
bc = 4x-3
bc = 4*7-3
bc = 25
====================
Cheers,
Stan H.
====================